OpenStudy (shiree19):

Evaluate the expression for s=4 and t=8. 2st^2 - s^2 I know the answer is 496 because that's was the correct answer in the book. How do you get this? Please explain. Thank you.

OpenStudy (shiree19):

\[2 \times 4 \times 8^{2} - 4^{2}; 8^{2} = 64; 64 \times 4 \times 2 = 512; 4^{2} = 16; 512 - 16 = 496\]

OpenStudy (shiree19):

Am I right?